Engineers want to design seats in commercial aircraft so that they are wide enough to fit 99​% of all males.​ (Accommodating 100% of males would require very wide seats that would be much too​ expensive.) Men have hip breadths that are normally distributed with a mean of 14.4 in. and a standard deviation of 0.9 in. Find Upper P 99. That​ is, find the hip breadth for men that separates the smallest 99​% from the largest 1​%.

The hip breadth for men that separates the smallest 99​% from the largest 1​% is Upper P 99= _______ in. ​(Round to one decimal place as​ needed.)
Answer #1

Given that,

mean = = 14.4

standard deviation = = 0.9

Using standard normal table,

P(Z < z) = 99%

= P(Z < z) = 0.99  

= P(Z <2.33 ) = 0.99

z =2.33 Using standard normal z table,

Using z-score formula  

x= z * +

x= 2.33*0.9+14.4

x= 16.497

x=16.5